OpenStudy (anonymous):

The following function is one-to-one. Find its inverse. Find the domain and range of f and f^-1 and Please show all of your work. f(x) = 8+x/x The following function is one-to-one. Find its inverse. Find the domain and range of f and f^-1 and Please show all of your work. f(x) = 8+x/x @Mathematics

OpenStudy (anonymous):

sorry disregard that

OpenStudy (anonymous):

so you have y=(8+x)/(x) correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

To find the inverse, switch x and y and solve for the new y.

OpenStudy (anonymous):

ok i got that part....but now its the domain and range of f and f^-1

OpenStudy (anonymous):

Easiest way is by using a graphing calculator, do you have one?

OpenStudy (anonymous):

yes. What do I put into though? lol

OpenStudy (anonymous):

first you try to graph the original function y=(8+x)/(x) and you do have a Ti calculator right?

OpenStudy (anonymous):

yes i have a Ti84 plus

OpenStudy (anonymous):

okay so once you graph that function go to 2nd GRAPH

OpenStudy (anonymous):

You will notice that there is going to be some ERROR signs on it, in this case it has one error sign right?

OpenStudy (anonymous):

well im putting in the original function and it tells me that there is an error in dim. Im going to y= then graph. What am I doing wrong?

OpenStudy (anonymous):

Deselect Plot1, Plot2, or Plot3 if any of them are enabled.

OpenStudy (anonymous):

Im not sure how to do that.

OpenStudy (anonymous):

wait i figured out that part

OpenStudy (anonymous):

on the y= tab there where you plug in the equation there is on the top where it says Plot1 Plot2 and Plot3 so if any of them are darkened you can just go up and click enter to deselect

OpenStudy (anonymous):

ok so the graph shows to go across the x-axis at -9, is that correct?!

OpenStudy (anonymous):

no, are you sure you have (8+x)/(x) plugged in just exactly how I wrote it?

OpenStudy (anonymous):

ohhh- haha- no i left off the parenthesis. y=8/x-1 ??

OpenStudy (anonymous):

but then how would i write the answer for the domain and range of f and f^-1?

OpenStudy (anonymous):

okay so you have f graphed right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now look at your graph

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

 Now it should look like this, does it?

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

okay so for domain you are looking at the x axis. So you determine where the graph starts and ends on whenever it is x. So as you can see this graph has two parts.

OpenStudy (anonymous):

when going to the left and one going to the right

OpenStudy (anonymous):

Now lets look at the one in the left side, in quadrant 3.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you are reading from left to right, where does the graph begin?

OpenStudy (anonymous):

for the first part?

OpenStudy (anonymous):

-7 ?

OpenStudy (anonymous):

 Look at the first part, it is coming from the negative direction